OpenStudy (anonymous):

Find the limit(if it exists) As lim as x goes to 10 of 10-x/x^2-100

myininaya (myininaya):

\[\lim_{x \rightarrow 10}\frac{-(x-10)}{(x-10)(x+10)}\]

myininaya (myininaya):

\[\lim_{x \rightarrow 10} \frac{-1}{x+10}\]

OpenStudy (anonymous):

so the answer is -1/20? or infinity?

OpenStudy (jamesj):

the function -1/(x+10) is continuous at x=10, so you can just substitute. Where would infinity come from now?

myininaya (myininaya):

x goes to -10 would be a different story

OpenStudy (anonymous):

Okay Thank You guys. Not sure why I said infinity